OpenStudy (anonymous):

Calculate Sin0 Its a right triangle We dont know the hypotonuse, but the Opposite is 6 and Adjcent is 7

OpenStudy (anonymous):



OpenStudy (cwrw238):

find hypotenuse using pythagoras = sqrt(6^2 + 7^2) then sin O = opposite / hypotenuse

OpenStudy (cwrw238):

you have opposite = 7 in your diagram but opposite = 6 in your question

OpenStudy (anonymous):

Drawing is correct

OpenStudy (cwrw238):

ok - so sin O = 7 / (sqrt(6^2 + 7^2)

OpenStudy (cwrw238):

can you calculate this?

OpenStudy (anonymous):

I got the answer! thank you! Does it work that way for all of he problems? For example: Calculate Cos0

OpenStudy (cwrw238):

yup cos O = adjacent / hypotenuse = 6 / hypotenuse have you heard of the memory aid SOH-CAH-TOA?

OpenStudy (anonymous):

I've heard of it.. but dont get it... So how do I set up the problem for Cos0 and Tan0

OpenStudy (phi):

you should use x if you cannot type \(\theta\). 0 looks like zero another way to do this is use tan x = opp/adj to find x, then do sin x or cos x

OpenStudy (cwrw238):

cos O is as i mentioned and tan O = opposite/adjacent = 7/6

OpenStudy (phi):

If you want more on this, you can watch http://www.khanacademy.org/math/trigonometry/v/basic-trigonometry

OpenStudy (anonymous):

so its 6/9.22 now how do I do COS

OpenStudy (phi):

cos x= 6/9.22= 0.65

OpenStudy (phi):

maybe the video will make this clear.